OpenStudy (anonymous):

What is the solution set for |2x – 8| 10?

OpenStudy (lilg132):

9 and -1

OpenStudy (anonymous):

u mean |2x – 8| = 10? right

OpenStudy (anonymous):

No \[\left| 2x - 8 \right| \ge 10\]

OpenStudy (lilg132):

oops then \[x \ge 9 \] and \[x \le\]

OpenStudy (nowhereman):

You should draw the graph of \[x ↦ |2x-8|\]

OpenStudy (lilg132):

\[x \le -1\]